OpenStudy (anonymous):

Find the indicated powers and express the answer in trigonometric form: [4(cos315+i sin315)]^2 Can someone help me please? Thanks ☺

OpenStudy (anonymous):

de Moivre's formula will help here: \[ [r(\cos(\theta)+i\sin(\theta))]^n = r^n\cos (n\theta) + i\sin (n\theta) \]